� The University of Electro–Communications � The University of Electro–Communications � The University of Electro–Communications^{� The University of Electro–Communications � The University of Electro–Communications � The University of Electro–Communications}^{� The University of Electro–Communications � The University of Electro–Communications � The University of Electro–Communications}^{Riichiro Saito  Riichiro Saito  Riichiro Saito}^{Riichiro Saito  Riichiro Saito  Riichiro Saito}^{Riichiro Saito  Riichiro Saito  Riichiro Saito}^{^{Associate Professor, Dr.}  ^{Associate Professor, Dr.}  ^{Associate Professor, Dr.}}^{^{Associate Professor, Dr.}  ^{Associate Professor, Dr.}  ^{Associate Professor, Dr.}}^{^{Associate Professor, Dr.}  ^{Associate Professor, Dr.}  ^{Associate Professor, Dr.}}^{^{Department of Electronic Engineering}  ^{Department of Electronic Engineering}  ^{Department of Electronic Engineering}}^{^{Department of Electronic Engineering}  ^{Department of Electronic Engineering}  ^{Department of Electronic Engineering}}^{^{Department of Electronic Engineering}  ^{Department of Electronic Engineering}  ^{Department of Electronic Engineering}}^{^{1-5-1, Chofugaoka Chofu, Tokyo, 182-8585, JAPAN}  ^{1-5-1, Chofugaoka Chofu, Tokyo, 182-8585, JAPAN}  ^{1-5-1, Chofugaoka Chofu, Tokyo, 182-8585, JAPAN}}^{^{1-5-1, Chofugaoka Chofu, Tokyo, 182-8585, JAPAN}  ^{1-5-1, Chofugaoka Chofu, Tokyo, 182-8585, JAPAN}  ^{1-5-1, Chofugaoka Chofu, Tokyo, 182-8585, JAPAN}}^{^{1-5-1, Chofugaoka Chofu, Tokyo, 182-8585, JAPAN}  ^{1-5-1, Chofugaoka Chofu, Tokyo, 182-8585, JAPAN}  ^{1-5-1, Chofugaoka Chofu, Tokyo, 182-8585, JAPAN}}^{^{e–mail: rsaito@ee.uec.ac.jp}  ^{e–mail: rsaito@ee.uec.ac.jp}  ^{e–mail: rsaito@ee.uec.ac.jp}}^{^{e–mail: rsaito@ee.uec.ac.jp}  ^{e–mail: rsaito@ee.uec.ac.jp}  ^{e–mail: rsaito@ee.uec.ac.jp}}^{^{e–mail: rsaito@ee.uec.ac.jp}  ^{e–mail: rsaito@ee.uec.ac.jp}  ^{e–mail: rsaito@ee.uec.ac.jp}}^{^{WWW: http://flex.ee.uec.ac.jp}  ^{WWW: http://flex.ee.uec.ac.jp}  ^{WWW: http://flex.ee.uec.ac.jp}}^{^{WWW: http://flex.ee.uec.ac.jp}  ^{WWW: http://flex.ee.uec.ac.jp}  ^{WWW: http://flex.ee.uec.ac.jp}}^{^{WWW: http://flex.ee.uec.ac.jp}  ^{WWW: http://flex.ee.uec.ac.jp}  ^{WWW: http://flex.ee.uec.ac.jp}}^{^{tel: +81-424-43-5148, fax: +81-424-43-5210}  ^{tel: +81-424-43-5148, fax: +81-424-43-5210}  ^{tel: +81-424-43-5148, fax: +81-424-43-5210}}^{^{tel: +81-424-43-5148, fax: +81-424-43-5210}  ^{tel: +81-424-43-5148, fax: +81-424-43-5210}  ^{tel: +81-424-43-5148, fax: +81-424-43-5210}}^{^{tel: +81-424-43-5148, fax: +81-424-43-5210}  ^{tel: +81-424-43-5148, fax: +81-424-43-5210}  ^{tel: +81-424-43-5148, fax: +81-424-43-5210}}